OpenStudy (anonymous):

Children with different ethnic backgrounds go to One World Charter School. Al of the students enrolled at the school speak English fluently. One-third of the students enrolled in the school can also speak Spanish fluently. One-third of the students who can speak Spanish fluently can also speak a third language. Whichof the following sentences correctly describes the fraction of students enrolled at One World Charter Schoolwho can speak three languages? A One-third multiplied by three equals one. B One-third multiplied by one equals one-third. C One-third divided by three equals one-ninth.

OpenStudy (anonymous):

D One-third divided by one-third equals one

OpenStudy (anonymous):

(1/3)(1/3) = fraction of students who speak three languages = (1/3)/3 = 1/9

OpenStudy (anonymous):

so is

OpenStudy (anonymous):

@douglaswinslowcooper so is

OpenStudy (anonymous):

C

OpenStudy (anonymous):

yeah thanks